Holton Company makes three products in a single facility. Data concerning these products follow:

Product
A B C
Selling price per unit $ 150.90 $ 76.30 $ 170.90
Direct materials $ 65.60 $ 42.50 $ 102.80
Direct labor $ 47.20 $ 13.50 $ 30.80
Variable manufacturing overhead $ 9.00 $ 4.50 $ 14.30
Variable selling cost per unit $ 24.40 $ 3.30 $ 8.80
Mixing minutes per unit 43.30 4.00 4.00
Monthly demand in units 3,000 1,000 2,000

The mixing machines are potentially the constraint in the production facility. A total of 14,000 minutes are available per month on these machines.

Direct labor is a variable cost in this company.

Required:

a. How many minutes of mixing machine time would be required to satisfy demand for all three products?

b. How much of each product should be produced to maximize net operating income?

c. Up to how much should the company be willing to pay for one additional hour of mixing machine time if the company has made the best use of the existing mixing machine capacity?
Answer #1

a) Required minutes = 3000*43.3+1000*4+2000*4 = 141900 minutes

b) Contribution margin per minute

A B C
Selling price 150.90 76.30 170.90
Variable cost 146.20 63.80 156.70
Contribution margin 4.70 12.50 14.20
Minute per unit 43.30 4 4
Contribution margin per minute 0.11 3.125 3.55

Optimum mix

Minute Unit
C 2000*4 = 8000 2000
B 4000 1000
A 2000 2000/43.30 = 46
Total 14000

c) Maximum price =